40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 2 10:22:37 2013, Rana, Gabriele, Jenne, Update, LSC, PRMI lock lock_prmi_sb.pdf
    Reply  Tue Apr 2 11:08:16 2013, Koji, Update, LSC, PRMI lock 
Message ID: 8387     Entry time: Tue Apr 2 10:22:37 2013     Reply to this: 8388
Author: Rana, Gabriele, Jenne 
Type: Update 
Category: LSC 
Subject: PRMI lock 

We locked the PRMI, this time really on the sidebands, using the two REFL55 signals.

Here are the parameters: triggering on POP22_I in at 140, out at 20. No normalization. MICH gain -0.15, PRCL gain 0.1

It seems that the lock is not very stable. It seems likely to come from some large angular motion of one of the mirrors. We'll need to calibrate the optical lever signals to understand which one is moving too much.

 

Attachment 1: lock_prmi_sb.pdf  240 kB  | Show | Show all
ELOG V3.1.3-